Need help ASAP with this question
Problem 2-4 Per-Share Earnings and Dividends [LO 2) 2 3
Answer #1

Please provide rating...

Answer 1) Computation of earning per share
earning per share = Net income/total number of share outstanding
a Sales 594000
Less expenses
Cost 264000
Depreciation expenses 66500
Interest expenses 33500
b Total expenses 364000
c=a-b Eearning before tax 230000
d=c*40% Tax @ 40% 92000
e=c-d Net income 138000
f Number of share outstanding 49000
g=e/f Earning per share          2.82
Answer 2) Computation of dividend per share
Dividend per share = Dividend paid/Number of share outstanding
a Total dividend paid = 40500
b Number of share outstanding 49000
c=a/b Dividend per share=          0.83